# Gaussian Building Mesh (GBM): Extract a Building’s 3D Mesh with Google Earth and Gaussian Splatting

### v2

@@ III-F Google Earth Studio
Google Earth Studio [1] is a web-based animation tool. With access to Google’s vast collection of 2D and 3D Earth data, ranging from large geological formations to individual buildings, Google Earth Studio provides a simple and efficient way to collect off-nadir images for the training of 3DGS models [47]. Google Earth Studio allows for the specification of a target of interest in terms of longitude and latitude coordinates, address, postal code, or location name. After selecting the target of interest, Google Earth Studio enables the specification of a camera path for which images of the target location are rendered. By selecting a circular camera path orbiting above and pointing towards the target of interest, Google Earth Studio allows for the extraction of a multi-view dataset of the building with 360-degree coverage, well suited for 3D reconstruction.
IV Methodology
@@
    Multi-view remote-sensing image collection: We leverage Google Earth Studio to collect multi-view images of a building of interest.
@@
    Building mask extraction: We extract multi-view consistent building masks for the building of interest in each of the images. We adapted existing GS2Mesh and SAM2 code to allow for text-based and click-based (point or box) user prompting to help SAM2 identify the building of interest in an initial image. We introduce re-prompting capabilities after mask propagation to address mask inconsistencies.
@@
    Building mask refinement: We refine SAM2 masks using morphological dilation, which slightly extends the mask outwards and fills in holes. We then simplify mask contours using the Ramer-Peuker-Douglas algorithm.
@@
    Gaussian Splatting: We train a 2DGS+ [5] model to learn building geometry and radiance.
@@
IV-A Google Earth Studio Dataset
We leverage Google Earth Studio to extract 7 scenes with buildings of interest that we wish to mesh. For each scene, we extract 31 frames/images over a 360°$\degree$ circular camera path centered around the building of interest. As part of Google Earth’s functionality, we tested and used a variety of methods to identify the buildings of interest, including address, postal code, geographic coordinates, and building name. Each scene’s camera information is provided in Table I. We rounded the camera tilt to the nearest half-degree. A camera tilt of 0°$\degree$ indicates the camera pointing straight down towards the ground, whereas a camera tilt of 90°$\degree$ indicates the camera pointing horizontally, parallel to the ground.
Location Camera Altitude (m) Camera Tilt (°$\degree$)
@@ IV-A Google Earth Studio Dataset
TABLE I: Camera altitude and tilt for Google Earth Studio scenes’ camera flight path
IV-B Mask Extraction and Mesh Refinement
SAM2 provides a rough mask of the building. We noticed the following problems while inspecting the SAM2 masks and the 3D mesh extraction that resulted from directly using those masks.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Tight-fitting mask, sometimes incomplete: SAM2 masks were often too tightly fitted onto buildings. This sometimes resulted in some outer parts of the buildings not being captured at certain viewpoints. This greatly affected 3DGS reconstruction quality.
  5. 5.
    Overly curved mask boundaries: SAM2 building masks have softer and rounder edges than expected. This is further exacerbated by using morphological dilation (see next subsection) to fill holes and extend building masks. We intuitively expect that, except for a few special cases, buildings are largely made of flat planar surfaces, which should result in masks with mostly straight-line-based boundaries.
To address these issues, we developed a mask refinement pipeline. We optionally morphologically erode the mask first, removing noise. Then, we grow the mask using morphological dilation to fill in any holes and expand incomplete masks. We then extract the mask’s contour using Canny’s algorithm and straighten/simplify the contour using the Ramer–Douglas–Peucker algorithm. The contour is then filled into a refined mask, addressing these issues.
IV-B1 User Re-prompting
Like GS2Mesh, by using the Grounded-SAM pipeline with SAM2, we extract object masks without further training via user prompting with natural language text (when combined with GroundingDINO), bounding boxes, or point-clicks. However, our preliminary testing showed that in a 360°$\degree$ video of a complex object, if the object self-occludes during rotation, the resulting masks can be inconsistent (see Figure 3). To solve this issue, we modified the GS2Mesh masking pipeline to add user re-prompting. We allow users to identify frames with errors and re-prompt the image at those frames. SAM2 re-propagates the user’s new prompt throughout all frames, effectively solving these mask inconsistencies.
@@ IV-B1 User Re-prompting
IV-B2 Morphological Dilation and Erosion
@@
where $I_{p}$ denotes the image $I$ shifted by pixel $p=(p_{x},p_{y})$. Roughly speaking, this grows the image by sliding the filter along the pixels of the original image. This operation can be performed iterative to further extend the image. At each step, the output image of the previous iteration is dilated again using the same filter. We apply two iterations of morphological dilation with a square filter of size (10,10).
@@
IV-B3 Canny Edge Extraction
@@
Canny edge extraction is a well-known classical algorithm for extracting lines using the image’s gradient. We briefly outline it here and refer readers to original paper for details [48]. We use a Gaussian filter of $\sigma=2$, a low threshold of 0, and a high threshold of 1 to extract a contour which we refine with the Ramer-Douglas-Peucker algorithm.
IV-B4 Ramer–Douglas–Peucker Contour Refinement
@@
The Ramer–Douglas–Peucker algorithm (RDP) [49] is an iterative contour simplification algorithm that recursively reduces the number of points in the contour.
This algorithm takes as input the list of points that define the curve and a distance threshold $\epsilon$.
  • If the curve has fewer than 3 points, it cannot be reduced further.
  • Find the point with the maximum perpendicular distance from the line formed by the first and last points of the curve.
  • If this distance is greater than $\epsilon$, it splits the curve at that point and recursively applies the algorithm to both parts.
  • If no point is farther than $\epsilon$, it replaces all the intermediate points with just the start and end points.
@@
We used $\epsilon=\frac{p}{500}$, as the ratio of the perimeter $p$ of the contour being simplified, allowing longer perimeter contours to retain more complexity. The resulting building contours became linear interpolations of approximately 20-40 points. We then fill the contour to create the refined mask which addresses the aforementioned issues.
IV-C 2DGS+
@@
We use a state-of-the-art implementation of Gaussian Splatting built on the foundations of 2DGS. This implementation, which we dubbed 2DGS+ [5], is an unpublished fork of the 2DGS repository and combines many recent advances in Gaussian Splatting, improving 2DGS with ideas from AbsGS [50], PixelGS [51], TrimGS [52], AtomGS [53], GaussianPro [54], and Taming-3DGS [55].
@@
In our experiments, we found that the Gaussian Splatting representations generated from 2DGS+ have fewer floaters (free-floating Gaussians in mid-air), are smoother, and produce better depth maps, resulting in significantly higher-quality meshes during TSDF fusion meshing (see section V-D). We tuned the hyperparameters for 3D mesh extraction quality (not novel view synthesis quality). We used the progressive training modality from GaussianPro, densifying the Gaussians until the 25,000th iteration. We used $\lambda_{depth}=0.2$ and $\lambda_{normal}=0.1$.
IV-D TSDF Fusion and Mesh extraction
@@
Like GS2Mesh and 2DGS, we used the Open3D implementation of the TSDF Fusion algorithm to convert depth maps into colored 3D meshes. We adapted the 2DGS mesh extraction module, adding a depth map smoothing step and tuning the hyperparameters. We used the 2DGS+ implementation [27] of 2DGS with its bounded mesh extraction module.
@@
V Results
Figure 5: Example of DLNR failure case. Left: 2D image of building of interest with background masked out. Right: DLNR stereo depth reconstruction from building with background masked out.
V-A Preliminary Experiments
Our preliminary experiments on 3D building mesh extraction from remote sensing images showed that the GS2Mesh 3D reconstruction pipeline, which uses DLNR [40], a pretrained neural-network-based stereo depth reconstruction and depth map fusion, only matched the quality of 2DGS depth map fusion in the best-case scenario. More often than not, it struggled in the 3D reconstruction of buildings due to two key issues: poor mask quality and poor stereo depth estimation. DLNR occasionally struggled with stereo depth reconstruction, likely because it was not trained on a remote sensing image dataset and has difficulty interpreting large-scale aerial scenes. This issue becomes particularly pronounced when stereo depth extraction is applied after masking, as the neural network loses the context of the background. In Figure 5, we observe that DLNR fails completely when performing depth extraction on a building with the background masked out. This motivated us to seek alternative solutions for the 3DGS-DLNR pipeline components in GS2Mesh.
Preliminary experiments with 2DGS+ showed that depth maps of certain complex buildings exhibited non-smoothness, resulting in poor-quality 3D meshes. We addressed this issue with the simple solution of smoothing the depth map using a Gaussian filter. An example is shown in Figure 4. After applying the filter, the holes in the zoomed-in region were almost completely filled. This particular issue arose due to poor visual coverage of the zoomed-in region in the initial training data.
Figure 6: Example of mask refinement with noise removal (top) and hole filling (bottom): Left: Raw SAM-2 Mask. Middle: Refined Mask. Right: masked Image. Regions of interest have been highlighted. Red indicates a region false positives or false negative pixels, blue indicates a region where the contour was refined and straightened.
V-B Mask Refinement
SAM2 masks sometimes included false-positive pixels away from the object of interest. In Figure 6 (top row), this issue is visible in the left-most subfigure. This is likely caused by false-positive segmentation due to the presence of a neighboring building’s rooftop with a similar shape and color in the original full image. Our mask refinement algorithm successfully removed the false-positive pixels and refined the mask contour.
@@ V-B Mask Refinement
V-C View Synthesis
Compared to GS2Mesh, we replaced the Gaussian Splatting module in the pipeline, transitioning from the original implementation of 3DGS to 2DGS+. As such, we compare the training set novel view synthesis results of 3DGS and 2DGS+ across our 7 scenes. We did not use the standard MipNeRF-360 convention of a training/testing split (leaving out one out of every eight images). Instead, we trained on the entirety of the Google Earth Studio footage, as we wanted smooth coverage of the 360°$\degree$ camera rotation for 3D reconstruction purposes. The novel view synthesis scores on the training set are provided in Table II.
On average, we found that 2DGS+ produced higher SSIM and LPIPS scores, whereas 3DGS achieved higher PSNR scores. We note that SSIM and LPIPS are typically better indicators of human visual perception than PSNR. Nonetheless, both models achieved novel view synthesis results that were nearly indistinguishable from ground truth images, with the exception of 3DGS struggling on the Canada Parliament scene, as shown in the top row of Figure 7 (see the central tower highlighted in the top-right zoom-in). This is also reflected in Table II, where the 3DGS scores for the Canada Parliament scene are significantly lower than for other scenes.
TABLE II: Comparison 2D Novel View Synthesis Metrics Between Gaussian Splatting Modules
@@ V-C View Synthesis
V-D 3D Building Mesh Extraction
Figure 8: Visualization of 3D colored mesh results. Left to Right: Canada Parliament; Dana Porter Library; ICON; Laurel Heights; CN Tower. Top row are our results, Bottom row are GS2Mesh results.
TABLE III: Comparison of 3D-SSIM Scores Between Our Pipeline and GS2Mesh
@@
As shown in Table III, our building extraction results are superior to GS2Mesh, with the exception of the Townhouse scene, where both methods scored similarly, within $\sim$0.0005 3D-SSIM. The results are somewhat skewed by the Canada Parliament scene (first column of Figure 8), where GS2Mesh 3D mesh reconstruction failed completely due to the significant presence of floaters during the training of its 3DGS module. This failure caused its DLNR depth reconstruction module to produce completely inaccurate depth maps. The poor mesh quality resulted in extremely low 3D-SSIM scores, as reflected in Table III. The resulting mesh was completely unrecognizable. The poor performance of the 3DGS module of GS2Mesh for this scene is also reflected in its PSNR score in Table II, which is approximately 10 dB lower than for other scenes. GS2Mesh also partially failed on the ICON scene and Dana Porter Library scene, where the lack of user re-prompting caused the masking to be inconsistent across the initial training images, resulting in missing geometric features and incomplete meshing.

### v3

@@ 3.6 Google Earth Studio
Google Earth Studio (Alphabet Inc., 2015-2024) is a web-based animation tool. With access to Google’s vast collection of 2D and 3D Earth data, ranging from large geological formations to individual buildings, Google Earth Studio provides a simple and efficient way to collect off-nadir images for the training of 3DGS models (Gao et al., 2024b). Google Earth Studio allows for the specification of a target of interest in terms of longitude and latitude coordinates, address, postal code, or location name. After selecting the target of interest, Google Earth Studio enables the specification of a camera path for which images of the target location are rendered. By selecting a circular camera path orbiting above and pointing towards the target of interest, Google Earth Studio allows for the extraction of a multi-view dataset of the building with 360-degree coverage, well suited for 3D reconstruction.
4 Methodology
@@
    Multi-view remote-sensing image collection: We leverage Google Earth Studio to collect multi-view images of a building of interest using its name, address, or postal code or coordinates.
@@
    Automated building mask extraction: We automatically extract multi-view consistent building masks for the building of interest in each image using system-level prompts.
@@
    Building mask refinement: We refine SAM2 masks using morphological dilation, which slightly extends the mask outwards and fills in holes. We then simplify mask contours using the Ramer-Peuker-Douglas algorithm. Optionally, we also allow for user-based re-prompting to correct possible errors from the automated process.
@@
    Gaussian Splatting: We train a 2DGS+ (Ye and GitHub-Contributors, 2024) model to learn building geometry and radiance.
@@
4.1 Google Earth Studio Dataset
We leverage Google Earth Studio to extract 7 scenes with buildings of interest that we wish to mesh. For each scene, we extract 31 frames/images over a 360°$\degree$ circular camera path centered around the building of interest. As part of Google Earth’s functionality, we tested and used a variety of methods to identify the buildings of interest, including address, postal code, geographic coordinates, and building name. Each scene’s camera information is provided in Table 1. We rounded the camera tilt to the nearest half-degree. A camera tilt of 0°$\degree$ indicates the camera pointing straight down towards the ground, whereas a camera tilt of 90°$\degree$ indicates the camera pointing horizontally, parallel to the ground.
4.2 Mask Extraction and Mesh Refinement
SAM2 provides a rough segmentation mask of the building. This process can be automated by providing system-level prompts to identify the building of interest. We noticed the following problems while inspecting the SAM2 segmentation masks and the 3D mesh extraction that resulted from directly using those masks.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Wrong building identification: When using automated system prompts in scenes where the building of interest is not a key landmark, (E.g. when multiple identical buildings are together), SAM2 at times segmented the wrong building.
  5. 5.
    Incomplete geometry: When using automated system-level prompting, at times, not the entire building of interest was segmented. Moreover, for complex buildings, SAM2 would at times lose part of the building in long image sequences.
To address these issues, we developed a mask refinement pipeline. We optionally morphologically erode the mask first, removing noise. Then, we grow the mask using morphological dilation to fill in any holes and expand incomplete masks. We then extract the mask’s contour using Canny’s algorithm and straighten/simplify the contour using the Ramer–Douglas–Peucker algorithm. The contour is then filled into a refined mask, addressing these issues. Our SAM2 image sequence-based building segmentation (mask generation) process is as follows.
4.2.1 Automated Mask Extraction
By using the Grounded-SAM pipeline with SAM2, we extract object masks without further training with natural language text. We use the system prompt "the central landmark or building of interest" to automatically extract the multi-view segmentation masks. Additionally, if a uniquely identifying proper noun (e.g. "The Parliament of Canada") was used to extract the multi-view images of the building from Google Earth Studio, we also automatically pass this to the Grounded-SAM pipeline as an additional system prompt.
4.2.2 User Re-prompting
To solve the aforementioned issue of incomplete geometry, we modified the masking pipeline to add optional user re-prompting. We allow users to identify frames with errors and re-prompt the image at those frames. SAM2 re-propagates the user’s new prompt throughout all frames, effectively solving these mask inconsistencies.
Table 1: Camera altitude and tilt for Google Earth Studio scenes’ camera flight path
Location Alt. (m) Tilt (°$\degree$)
@@ 4.2.2 User Re-prompting
4.2.3 Morphological Dilation and Erosion
@@
where $I_{p}$ denotes the image $I$ shifted by pixel $p=(p_{x},p_{y})$. Roughly speaking, this grows the image by sliding the filter along the pixels of the original image. This operation can be performed iteratively to further extend the image. At each step, the output image of the previous iteration is dilated again using the same filter.
@@
4.2.4 Canny Edge Extraction
@@
Canny edge extraction is a well-known classical algorithm for extracting lines using the image’s gradient. The Canny algorithm is well-known in computer vision. As such, we refer readers to the original paper for details (Canny, 1986).
4.2.5 Ramer–Douglas–Peucker Contour Refinement
@@
The Ramer–Douglas–Peucker algorithm (RDP) (Ramer, 1972) is an iterative contour simplification algorithm that recursively reduces the number of points in the contour. The algorithm is as follows:
Given: A curve defined by a sequence of ordered points: $P=\{P_{1},P_{2},\ldots,P_{n}\},\quad\text{where }P_{i}=(x_{i},y_{i})$ (2) and a threshold parameter: $\epsilon>0$ (3) Algorithm: 1. Start with endpoints $P_{1}$ and $P_{n}$. Define the line segment: $\overline{P_{1}P_{n}}$ 2. For each point $P_{i}=(x_{i},y_{i})$ with $i\in[2,n-1]$, compute the perpendicular distance $d_{i}$ to the line $\overline{P_{1}P_{n}}$: $d_{i}=\frac{|(x_{n}-x_{1})(y_{1}-y_{i})-(x_{1}-x_{i})(y_{n}-y_{1})|}{\sqrt{(x_{n}-x_{1})^{2}+(y_{n}-y_{1})^{2}}}$ (4) 3. Determine the maximum distance: $d_{\text{max}}=\max_{i=2,\ldots,n-1}d_{i}$ (5) 4. If $d_{\text{max}}>\varepsilon$, let $P_{k}$ be the point at which the maximum distance occurs. Recursively apply the algorithm to the subsets: $\{P_{1},P_{2},\ldots,P_{k}\}\quad\text{and}\quad\{P_{k},P_{k+1},\ldots,P_{n}\}$ (6) 5. If $d_{\text{max}}\leq\varepsilon$, approximate the entire set by the line segment $\overline{P_{1}P_{n}}$, and keep only $P_{1}$ and $P_{n}$. Output: A reduced set of points $P^{\prime}\subseteq P$, forming a piecewise linear approximation within tolerance $\epsilon$.
@@
We used $\epsilon=\frac{p}{500}$, as the ratio of the perimeter $p$ of the contour being simplified, allowing longer perimeter contours to retain more complexity. The resulting building contours became linear interpolations of approximately 20-40 points. We then filled the contour to create the refined mask, which addresses the aforementioned issues. We tested both our own implementation and an OpenCV implementation and found no discernible difference.
4.3 2DGS+
@@
We use a state-of-the-art implementation of Gaussian Splatting built on the foundations of 2DGS. This implementation, which we dubbed 2DGS+ (Ye and GitHub-Contributors, 2024), is an unpublished fork of the 2DGS repository and combines many recent advances in Gaussian Splatting, improving 2DGS with ideas from AbsGS (Ye et al., 2024), PixelGS (Zhang et al., 2024), TrimGS (Fan et al., 2024), AtomGS (Liu et al., 2024), GaussianPro (Cheng et al., 2024), and Taming-3DGS (Mallick et al., 2024).
@@
In our experiments, we found that the Gaussian Splatting representations generated from 2DGS+ have fewer floaters (free-floating Gaussians in mid-air), are smoother, and produce better depth maps, resulting in significantly higher-quality meshes during TSDF fusion meshing (see section 5.4). We tuned the hyperparameters for 3D mesh extraction quality (not novel view synthesis quality). We used the progressive training modality from GaussianPro, densifying the Gaussians until the 25,000th iteration. We used $\lambda_{depth}=0.2$ and $\lambda_{normal}=0.1$.
4.4 TSDF Fusion and Mesh extraction
@@
Like GS2Mesh and 2DGS, we used the Open3D implementation of the TSDF Fusion algorithm to convert depth maps into colored 3D meshes. We adapted the 2DGS mesh extraction module, adding a depth map smoothing step and tuning the hyperparameters. We used the 2DGS+ implementation (Huang et al., 2024) of 2DGS with its bounded mesh extraction module.
@@
5 Results
Although there are multiple Gaussian Splatting-based methods for 3D mesh extraction of entire scenes, we found GS2Mesh to be the only method comparable in its application: mesh extraction for single objects from text-based or click-based prompts with optional automation. Additionally, we used it as a starting point to develop our framework. As such, we focus our comparisons against it.
Figure 5: Example of DLNR failure case. Left: 2D image of the building of interest with background masked out. Right: DLNR stereo depth reconstruction from building with background masked out.
5.1 Preliminary Experiments
Our preliminary experiments on 3D building mesh extraction from remote sensing images showed that the GS2Mesh 3D reconstruction pipeline, which uses DLNR (Zhao et al., 2023), a pretrained neural-network-based stereo depth reconstruction and depth map fusion, only matched the quality of 2DGS depth map fusion in the best-case scenario. More often than not, it struggled in the 3D reconstruction of buildings due to two key issues: poor mask quality and poor stereo depth estimation. DLNR occasionally struggled with stereo depth reconstruction, likely because it was not trained on a remote sensing image dataset and had difficulty interpreting large-scale aerial scenes. This issue becomes particularly pronounced when stereo depth extraction is applied after masking, as the neural network loses the context of the background. In Figure 5, we observe that DLNR fails completely when performing depth extraction on a building with the background masked out. This motivated us to seek alternative solutions for the 3DGS-DLNR pipeline components in GS2Mesh.
Additional preliminary experiments with 2DGS+ showed that depth maps of certain complex buildings exhibited non-smoothness, resulting in poor-quality 3D meshes. We addressed this issue with the simple solution of smoothing the depth map using a Gaussian filter. An example is shown in Figure 4. After applying the filter, the holes in the zoomed-in region were almost completely filled. This particular issue arose due to poor visual coverage of the zoomed-in region in the initial training data.
Figure 6: Example of mask refinement with noise removal (top) and hole filling (bottom): Left: Raw SAM-2 Mask. Middle: Refined Mask. Right: masked Image. Regions of interest have been highlighted. Red indicates a region with false positives or false negative pixels, blue indicates a region where the contour was refined and straightened.
5.2 Mask Refinement
SAM2 masks sometimes included false-positive pixels away from the object of interest. This issue is visible in the left-most subfigure of Figure 6 (top row). This is likely caused by false-positive segmentation due to the presence of a neighboring building’s rooftop with a similar shape and color in the original full image. Our mask refinement algorithm successfully removed the false-positive pixels and refined the mask contour.
@@ 5.2 Mask Refinement
Table 2: Comparison 2D Novel View Synthesis Metrics Between Gaussian Splatting Modules
@@
Table 3: Comparison of 3D-SSIM Scores Between Our Pipeline and GS2Mesh
@@
5.3 View Synthesis
Compared to GS2Mesh, we replaced the Gaussian Splatting module in the pipeline, transitioning from the original implementation of 3DGS to 2DGS+. As such, we compare the training set novel view synthesis results of 3DGS and 2DGS+ across our 7 scenes. We did not use the standard MipNeRF-360 convention of a training/testing split (leaving out one out of every eight images). Instead, we trained on the entirety of the Google Earth Studio footage, as we wanted smooth coverage of the 360°$\degree$ camera rotation for 3D reconstruction purposes. The novel view synthesis scores on the training set are provided in Table 2.
On average, we found that 2DGS+ produced higher SSIM and LPIPS scores, whereas 3DGS achieved higher PSNR scores. We note that SSIM and LPIPS are typically better indicators of human visual perception than PSNR. Nonetheless, both models achieved novel view synthesis results that were nearly indistinguishable from ground truth images, with the exception of 3DGS struggling on the Canada Parliament scene, as shown in the top row of Figure 7 (see the central tower highlighted in the top-right zoom-in). This is also reflected in Table 2, where the 3DGS scores for the Canada Parliament scene are significantly lower than for other scenes.
5.4 3D Building Mesh Extraction
Figure 8: Visualization of 3D colored mesh results, showcasing imperfections in both methods. Left to Right: Canada Parliament; Dana Porter Library; ICON; Laurel Heights; CN Tower. Top row are our results, Bottom row are GS2Mesh results.
As shown in Table 3, our building extraction results are superior to GS2Mesh, with the exception of the Townhouse scene, where both methods scored similarly, within $\sim$0.0005 3D-SSIM. The results are somewhat skewed by the Canada Parliament scene (first column of Figure 8), where GS2Mesh 3D mesh reconstruction failed completely due to the significant presence of floaters during the training of its 3DGS module. This failure caused its DLNR depth reconstruction module to produce completely inaccurate depth maps. The poor mesh quality resulted in extremely low 3D-SSIM scores, as reflected in Table 3. The resulting mesh was completely unrecognizable. The poor performance of the 3DGS module of GS2Mesh for this scene is also reflected in its PSNR score in Table 2, which is approximately 10 dB lower than for other scenes. GS2Mesh also partially failed on the ICON scene and Dana Porter Library scene, where the lack of user re-prompting caused the masking to be inconsistent across the initial training images, resulting in missing geometric features and incomplete meshing.